OpenStudy (anonymous):

The usual toll charge to use the Bingham tunnel is 50 cents. If you purchase a special sticker for $5.50, the toll is only 35 cents. At least how many trips through the tunnel are needed before the sticker costs less than paying for each trip separately?

OpenStudy (anonymous):

trips taken at usual toll = 0.50 * x = 0.50x trips taken with sticker = 5.50 + 0.35x You want to know when they're equal to each other. 0.50x = 5.50 + 0.35x 0.15x = 5.50 x = 5.50/0.15 x = 36 2/3 Check: 0.50x = 0.50 * 36 2/3 = 18 1/3 5.50 + 0.35x = 5.50 + (0.35 * 36 2/3) = 5.50 + 12.8333333 =18 1/3 True